OpenStudy (anonymous):

Complete the Next Step: A partial solution to the equation 16 + 10x = 14x − 4 is given below. Choose the step that is missing from the solution. 16 + 10x = 14x − 4 16 = 4x − 4 _____________ _____________ 5 = x

OpenStudy (homeworksucks):

adding 4 to both sides, then dividing both sides by 4

OpenStudy (anonymous):

16+10x = 14x-4 16+4 = 14x-10x 20 = 4x x = 5

OpenStudy (homeworksucks):

what he said